Zelynross:

Multiply the complex numbers: (2 + i)(3 - 6i) *

Zelynross:

@XioGonz

XioGonz:

You need to know the Formula The formula is :\[(a+bi) • (x+yi) = ax+by•i^2+(ay+bx)i \] Now, since \[ i^2 = -1 \] it can be written as :\[ax - by + (ay + bx) i \] And\[ ( 2 + i) • ( 3 - 6i) =\] Hence \[( 12 - 9i) \]

Zelynross:

preciate it

XioGonz:

Np